OpenStudy (anonymous):

For what values of x is the following expression undefined? (3x – 2)(5x – 4) / (9x – 6)(7x – 1) Show how you obtained your answer. Why is the answer important?

OpenStudy (amistre64):

when the underside goes zero it is undefined

OpenStudy (amistre64):

a*b = 0 when a=0 or b=0, or both = 0

OpenStudy (anonymous):

I have this as my steps 1. (3x-2) (5x-4) ----------------- (9x-6) (7x-1) 2. Factor out 3 from each term in polynomial. (3x-2) (5x-4) ----------------------------- (3(3x)+3(-2))(7x-1) 3. Factor out the 3 from 9x-6 (3x-2)(5x-4) ------------------------- (3(3x-2))(7x-1) 4. 3(3x-2) (7x-1)=0 This would make the fraction undefined.

OpenStudy (amistre64):

good, but what values of x make it go undefined?

OpenStudy (anonymous):

and then 6.Solve the equation to find where the original expression is undefined X=2/3,1/7 7.The answer is important because if the denominator is equaled to zero, the expression cannot be solved. I cant figure out number 5

OpenStudy (amistre64):

it looks good, but what is 5 asking for?

OpenStudy (anonymous):

like step i had it all right just my teacher wants me to go back and make the step's clear and i dont know what to put

OpenStudy (amistre64):

i dont know what your teacher would want either. to be clearer might mean to define your actions by stated axioms and thrms

OpenStudy (anonymous):

1. (3x-2) (5x-4) ----------------- (9x-6) (7x-1) 2. Factor out 3 from each term in polynomial. (3x-2) (5x-4) ----------------------------- (3(3x)+3(-2))(7x-1) 3. Factor out the 3 from 9x-6 (3x-2)(5x-4) ------------------------- (3(3x-2))(7x-1) 4. Then remove the parentheses around the 3. 3(3x-2) (7x-1)=0 This would make the fraction undefined. 5.Solve the equation to find where the original expression is undefined X=2/3,1/7 6.The answer is important because if the denominator is equaled to zero, the expression cannot be solved. She says please show a little more detail in how you got from step 4 to step 5?.. I cant think of any other way

OpenStudy (amistre64):

i got it

OpenStudy (amistre64):

split step 4 into its component parts and solve each part for 0

OpenStudy (amistre64):

(3x-2) (7x-1)=0 ; since a*b = 0 when .... 3x-2=0 7x-1=0

OpenStudy (anonymous):

okay so put that where step number 4 is?

OpenStudy (amistre64):

include that into step 4, its transitioning from 4 to 5

OpenStudy (amistre64):

its the part that was done in your head that they wants on paper

OpenStudy (anonymous):

ur the best lol

OpenStudy (amistre64):

;)